OpenStudy (anonymous):

If you solve a trig equation and you get the solutions shown below, find all solutions on the interval [-2pi, 2pi)

OpenStudy (anonymous):

\[[\frac{ \pi}{ 4 }+2\pi n, \frac{ \pi }{ 3 }+ 2\pi n, n \in \mathbb{Z}^+ ]\]

OpenStudy (freckles):

that will give the solutions between [0,2pi] if you set n=0 that is consider the first rotation if you had wanted the solutions in [2pi,4pi] you set n=1 can you make a guess as to what you what set n to in order to get the solutions in [-2pi,0]?

OpenStudy (anonymous):

n= -1 ?

OpenStudy (freckles):

yes

OpenStudy (freckles):

so to get the solutions in [-2pi,0] set n=-1 to get the solutions in [0,2pi] set n=0 this will give you all the solutions in [-2pi,2pi]

OpenStudy (anonymous):

ok so use the solutions given above and set n to -1 and 0 to find the other solutions?

OpenStudy (freckles):

you will replace all the n's with -1 then you will replace all the n's with 0

OpenStudy (freckles):

the weird thing is Z^+ means positive integers only

OpenStudy (freckles):

So that is odd and weird since to find the solutions in [-2pi,2pi] we need non positive n's

OpenStudy (anonymous):

o ok, it was supposed to be all all solutions, not only positive. my bad

OpenStudy (anonymous):

anyways, thanks for showing me the way how to do it, bud. I was confused, reading the question and was not sure what to do.

OpenStudy (freckles):

np